OpenStudy (anonymous):

Which of the following exponential functions goes through the points (1, 12) and (2, 36)?

OpenStudy (victoriasushchik):

if you plug in x = 1 we have 4(3)^1 = 12 if you plug in x = 2 we have y = 4(3)^2 = 4*9 = 36

OpenStudy (anonymous):

Thanks!

OpenStudy (victoriasushchik):

your welcome Hon!